OpenStudy (matt6288):

Given a circle with a diameter of 4, what is the circumference?

OpenStudy (saifoo.khan):

\[Radius = \frac{diameter}{2} = \frac 42 = 2\]

OpenStudy (anonymous):

pi * diameter = circumference pi * 4 = given circle's circumference

OpenStudy (saifoo.khan):

\[Circ = 2 \pi r\] \[2 \pi 2\]

OpenStudy (saifoo.khan):

Use a calc now! :)

OpenStudy (anonymous):

it is 12.56 i think

OpenStudy (matt6288):

A. 3 pie B.4 pie c.6pie D.8pie

OpenStudy (anonymous):

B

OpenStudy (anonymous):

its b for sure